# Thinking with Images for Multimodal Reasoning: Foundations, Methods, and Future Frontiers

### v2

@@ 8 Future Directions
“Thinking with Images” marks a paradigm shift in artificial intelligence; yet, we are only at the cusp of realizing its full potential. Looking forward, this section explores the pivotal research avenues that will propel the field into its next era, focusing on the innovations required to unlock more efficient, secure, and profoundly capable forms of visual cognition.
@@ 8.1 Computational and Cognitive Efficiency
A critical frontier for “Thinking with Images” is the enhancement of its operational efficiency (Ning et al., 2024; Zhou et al., 2024c). While current paradigms have unlocked unprecedented capabilities, they often depend on sequential visual execution, such as invoking external tools or generating intermediate images. This approach introduces significant computational overhead and latency, rendering these powerful models impractical for real-time applications (Liu et al., 2024b). Future research must therefore develop methods to achieve deep visual cognition without such costly execution. This involves compressing multi-step deliberation into a more compact form. Models could learn computational shortcuts or anticipate the outcome of a reasoning chain without rendering every intermediate visual state (Lee et al., 2025).
Beyond computational speed, a more profound challenge lies in achieving cognitive efficiency, which concerns the model’s ability to judiciously allocate its reasoning resources (Qu et al., 2025; Huang et al., 2025c). A truly intelligent system should not apply its most expensive visual thinking capabilities to every problem (Wang et al., 2025m). The next generation of models must learn when a simple textual inference is sufficient versus when a deeper, multimodal analysis is necessary. This requires a metacognitive framework to assess task complexity and dynamically select the most economical reasoning pathway. Future work could explore reward mechanisms that penalize unnecessary visual operations to encourage an intrinsic sense of inferential parsimony. The goal is to transition from models that follow a rigid procedure to ones that can flexibly interleave fast textual processing with more deliberate visual analysis as the context demands.
Achieving this dual efficiency will also likely require architectural innovations (Deng et al., 2025a; Chen et al., 2025b). Instead of treating visual operations as external functions, future architectures could feature tightly integrated, lightweight modules. These native modules could be designed for specific cognitive functions, such as visual search or element comparison. By operating directly within the model’s latent space, they would bypass the overhead of pixel-level manipulation. The objective is to cultivate models that not only “Thinking with Images” but do so with an innate sense of purpose and economy, making sophisticated visual reasoning a scalable and accessible reality.
@@ 8.2 Safe and Trustworthy Visual Cognition
Beyond performance, the responsible development of models that “Thinking with Images” introduces unique ethical vulnerabilities (Andriushchenko et al., 2024; Qu et al., 2024b). A primary danger lies in the model’s capacity for internal generation, which can be weaponized to construct highly persuasive visual arguments (Nguyen et al., 2022). A malicious actor could guide the model to produce a series of fabricated images that appear as logical evidence for a false conclusion. This capability moves beyond simple deepfakes toward automated and scalable disinformation campaigns. The challenge is therefore not just to prevent the creation of a single harmful image, but to ensure the entire cognitive process cannot be co-opted to rationalize a malicious narrative, thereby safeguarding information integrity (Bai et al., 2022).
The security of these models is also challenged by their complex cognitive architecture. The attack surface expands from the initial input to the entire internal reasoning loop (Liu et al., 2024c, e, f). An adversary could compromise the external tools that the model relies upon, poisoning its analysis at the source. More subtly, an attacker could design inputs that manipulate the model’s autonomous choices. This might trick the model into selecting a flawed operation or focusing on a distracting detail, effectively derailing its thought process. The security threat thus evolves from generating an incorrect output to corrupting the mechanisms of cognition itself. This shift demands new defenses that can protect the integrity of the model’s decisions, its tools, and its internal visual imagination.
Finally, the opaque nature of this proactive thinking process presents an obstacle to trustworthiness. Algorithmic bias can become embedded not just in the final output, but within the model’s internal operations (Howard et al., 2024). A system might learn to apply greater scrutiny or generate stereotypical imagery when reasoning about certain demographics. This could create a self-reinforcing cycle of prejudice that is difficult to detect. Making these systems transparent requires more than explaining a result (Aflalo et al., 2022). It demands methods to visualize and audit the model’s internal dialogue of tool use and image generation. Without such interpretability, diagnosing biases or understanding failures becomes impossible. This limitation hinders our ability to build models that are not only capable but also demonstrably fair and aligned with human values (Chang et al., 2024).
@@ 8.3 Novel Benchmarks and Evaluation Methodologies
A significant challenge for the “Thinking with Images” paradigm is the inadequacy of current evaluation benchmarks (Li et al., 2024f). Prevailing datasets, designed for tasks such as Visual Question Answering, primarily measure final output accuracy. They often fail to scrutinize the underlying reasoning process. This limitation creates a critical gap, as models can achieve high scores by exploiting dataset biases or employing textual shortcuts, without genuinely integrating visual evidence into their cognitive workflow (Chen et al., 2024d). Consequently, we lack robust methods to differentiate between authentic visual reasoning and sophisticated pattern matching (Wang et al., 2024h). This ambiguity hinders our ability to gauge progress in active visual cognition and to diagnose the failure modes of models that appear to reason but do not truly understand.
Future research could prioritize the development of benchmarks engineered to probe the model’s reasoning trajectory (Lotfi et al., 2024). These benchmarks should move beyond static query-response pairs to feature multi-step problems. In such tasks, visual information must be indispensable for each stage of inference, not merely contextual. For instance, tasks could involve visual abductive reasoning, where a model infers the most plausible cause of a scene. Other challenges might require synthesizing information from multiple, disparate image regions. The design of such tasks would necessitate a deep integration of visual perception and logical deduction, ensuring that success is contingent upon the model’s ability to actively think with the image (Wu et al., 2024h).
The creation of new benchmarks must be accompanied by innovative evaluation methodologies. Moving beyond simple accuracy scores, future metrics must assess the quality and faithfulness of the reasoning process itself (Song et al., 2025). This requires developing systems to evaluate how well each step in a thought process is grounded in specific visual evidence, thereby penalizing hallucinated or unfaithful reasoning. Furthermore, these methodologies should evaluate the logical coherence and causal necessity of the entire reasoning chain. Establishing such process-oriented evaluation protocols is a foundational step toward building more transparent and reliable multimodal systems capable of complex, image-centric thought.
@@ 8.4 Thinking with Audio, Video, and the World
The ultimate frontier of this research culminates in “Thinking with the world”, where the paradigm shifts from detached observation to embodied agency (Brohan et al., 2023). Here, reasoning becomes an integral component of a continuous perception-action loop. The model’s thoughts directly inform its physical actions, and the sensory consequences provide immediate feedback that grounds its understanding (Fu et al., 2024b). Its generative power evolves into a mechanism for mental rehearsal, allowing the agent to simulate action sequences before committing to a physical choice. This process grounds abstract concepts in tangible reality, defining an object not by its appearance but by its physical affordances in a planned task (Liu et al., 2024g).
@@ 8.5 Open Questions and Outlook
Rather than being mutually exclusive, the “Thinking with Images” paradigm and an agent framework share a deeply synergistic and nested relationship. The most effective way to conceptualize this is to view the “Thinking with Images” paradigm as the cognitive engine within the agent’s machine. An agent provides the essential architecture for worldly interaction, equipped with perception sensors, action effectors, and memory systems. The “Thinking with Images” paradigm provides the sophisticated reasoning power that makes the agent’s actions intelligent rather than merely reactive. This synergy becomes tangible when we consider how the stages outlined in this survey empower an agent’s capabilities. An agent leverages Stage 1 by learning to orchestrate external visual tools. It evolves by using Stage 2 to programmatically analyze visual data. Ultimately, a truly advanced agent embodies Stage 3, using intrinsic imagination to simulate the consequences of potential actions. Therefore, the research presented in this survey is not an alternative to agent-based AI but is instead a foundational step toward building autonomous systems that possess a deeper, more human-like capacity for visual understanding and intelligent action.
@@ 9 Conclusion
The journey towards true visual cognition is still in its early stages. Significant challenges in efficiency, robustness, and generalization must still be overcome. The path forward requires developing more sophisticated world models, exploring abstract visual representations beyond pixels, and building unified frameworks for human-AI collaboration. Ultimately, the research presented here marks a foundational step toward creating AI agents that not only see the world, but reason within it as well, equipped with a mind’s eye that is integral to their intelligence.
Appendix A Appendix
A.1 Expanded Benchmark Details
The following tables provide a comprehensive overview of the benchmarks discussed in Section 6.
Benchmark Characteristics Scenarios Eval Metrics
Setup
Mathematical Reasoning
GeoQA+ [Cao and Xiao, 2022] Geometry problems with complex visual elements P R G Auto
UniGeo [Chen et al., 2022a] Unified text-diagram geometric problem solving P R G Auto
TrustGeoGen [Fu et al., 2025b] Generating trustworthy geometric diagrams P G G Auto
GeoEval [Zhang et al., 2024f] Geometric reasoning from textual descriptions P R M G Auto
VisioMath [Li et al., 2025b] Math word problems with visual diagram answers P R M Auto
MV-MATH [Wang et al., 2025f] Math reasoning in multi-visual contexts P R M G Auto
MathVista [Lu et al., 2024] Large-scale, multi-task visual math reasoning P R M G Hybrid
MM-MATH [Sun et al., 2024c] Middle school math with reasoning process evaluation P R G Hybrid
MathVision [Wang et al., 2024c] Competition-level multimodal math problems P R G Auto
MathVerse [Zhang et al., 2024d] Multi-version diagrammatic reasoning for robustness P R G Hybrid
OlympiadBench [He et al., 2024b] Olympiad-level math and physics problems P R G Hybrid
LogicVista [Xiao et al., 2024] Visio-linguistic logical reasoning P R M Auto
DynaMath [Zou et al., 2025] Dynamic benchmark for robustness evaluation P R G Auto
VisOnlyQA [Kamoi et al., 2024] Visual-only mathematical problems P M Auto
MathGlance [Sun et al., 2025] Quick assessment of mathematical capabilities P R M Auto
MathExplain [Park et al., 2025] Generating natural language math explanations R G LLM
Polymath [Gupta et al., 2024a] Cognitive multimodal challenges for high-level mathematical reasoning P R G Auto
STEM
ScienceQA [Lu et al., 2022] Multimodal science questions with explanations P R M Auto
MMIE [Xia et al., 2025b] Comprehensive multimodal intelligence evaluation P R M G Hybrid
EMMA [Hao et al., 2025] Educational material analysis from textbooks P R M G Auto
MDK12-Bench [Zhou et al., 2025b] K-12 multimodal science education P R M G Auto
Isobench [Fu et al., 2024a] Scientific reasoning with isomorphic problems P R M G Auto
CSVQA [Jian et al., 2025] Visual question answering on scientific charts P R M G Auto
SWE-bench MM [Yang et al., 2025b] Multimodal software engineering tasks P R G G Auto
MMMU [Yue et al., 2024a] Massive multimodal university-level subjects P R M Auto
MMMU-Pro [Yue et al., 2024b] Professional-level expert reasoning P R G M Auto
SFE [Zhou et al., 2025a] Expert-level scientific reasoning P R G Auto
MMSciBench [Ye et al., 2025] Multimodal scientific reasoning P R M G Auto
PhysUniBench [Wang et al., 2025j] Undergraduate-level multimodal physics problems P R G M G Auto
Puzzles and Games
INGVP [Zhang et al., 2024e] Spatial imagination & rule-following puzzles P R M Auto
LVLM-Playground [Wang et al., 2025l] Multi-step decision making in classic video games P R G Auto
GameQA [Tong et al., 2025a] Game logic and state understanding P R M Auto
BALROG [Paglieri et al., 2024] Rule comprehension in text-based games P R G Hybrid
VGRP-Bench [Ren et al., 2025] Grid-based puzzles with logical reasoning P R G Auto
PuzzleBench [Zhang et al., 2025c] Visual QA based on puzzles P R G Auto
VideoGameBench [Zhang et al., 2025g] Understanding dynamic video game data P R G Auto
PhysGame [Cao et al., 2024] Physical reasoning & error detection in games P R M Auto
VCBench [Li et al., 2024b] Visual commonsense, abstraction, and control P R G LLM
VSP [Wu et al., 2024f] Perception and reasoning in spatial planning P R G G Auto
EnigmaEVAL [Wang et al., 2025k] Multi-step reasoning on unstructured puzzles from competitions P R G G Auto
PuzzleWorld [Li et al., 2025h] Large-scale benchmark for creative open-ended puzzle solving P R G G Hybrid
Table 5: Benchmarks assessing the “Thinking with Images” paradigm across Mathematical Reasoning, STEM, and Puzzles and Games. The columns detail each benchmark’s Characteristics; the core cognitive Scenarios tested (P: Active Perception, R: Visual Reasoning, G: Generative Planning); the Eval Setup (M: multi-choice, G: generation); and the evaluation Metrics (Auto: automatic metrics, LLM: LLM-as-a-Judge, Hybrid: a combination of both).
Benchmark Characteristics Scenarios Eval Metrics
Setup
Perception Reasoning
V*Bench [Wu and Xie, 2023] Fine-grained Perception on High-Resolution Images P R M Auto
BLINK [Huang et al., 2023] Fine-grained Visual Perception P M Hybrid
HR-Bench [Wang et al., 2024e] Fine-grained Reasoning on High-Resolution Images P R M G Auto
Verify [Kumar et al., 2024] Visual Explanation and Multi-step Logical Reasoning P R G LLM
POPE [Li et al., 2023a] Object-level Hallucination Detection P R M Auto
H-POPE [Li et al., 2024e] Hierarchical Object Hallucination Detection P R M Auto
CogBench [Zheng et al., 2024c] Assessing Cognitive Abilities of LMMs P R G LLM
RISE-Bench [Gupta et al., 2024b] Evaluating Reasoning-Informed Visual Editing R G G LLM
VBench [Wu et al., 2023b] Comprehensive Video Generation Evaluation G G Hybrid
Code Generation
MMCod [Li et al., 2024c] Visual Content-Grounded Code Generation P R G G Auto
Plot2Code [Wu et al., 2025b] Mathematical Plot-to-Code Generation P G G Hybrid
Design2Code [Si et al., 2025] Design Layout-to-Webpage Generation P G G Auto
Sketch2Code [Li et al., 2025i] Hand-drawn Sketch-to-Code Generation P G G Hybrid
WebSight [Laurençon et al., 2024] Web Screenshot-to-Code generation P G G Hybrid
Chart and Table Reasoning
ChartX [Xia et al., 2024b] Chart Type Recognition and Data Extraction P R G M G Hybrid
ChartBench [Xu et al., 2023] Chart Information Extraction and Trend Analysis P R M G Hybrid
CharXiv [Wang et al., 2024d] Scientific Chart Understanding and QA P R M G Hybrid
ChartSumm [Tang et al., 2025] Chart Summarization P R G G LLM
MultiChartQA [Zhu et al., 2025c] Multi-chart Question Answering and Reasoning P R M G Hybrid
TableBench [Wu et al., 2025c] Tabular Data Analysis and QA P R G Auto
M3SciQA [Li et al., 2024d] Multimodal Scientific Chart and Table QA P R M Hybrid
ChartMimic [Shi et al., 2024b] Visually-grounded Chart-to-Code Generation P G G Auto
Ocr-Reasoning [Huang et al., 2025b] Reasoning on OCR-rich Images P R G Hybrid
Real-world Applications
MULTI [Zhu et al., 2024] Diverse Real-World Multimodal Tasks P R G Hybrid
M3GIA [Zhang et al., 2023c] Cross-Disciplinary Multimodal Exam QA P R M Auto
ClockQA [Saxena et al., 2025] Time Reading and Reasoning on Clock Images P R M G Hybrid
CalendarQA [Saxena et al., 2025] Date Understanding and Reasoning on Calendars P R M G Hybrid
VisNumBench [Weng et al., 2025b] Visual Numerical Reasoning and Counting P R M G Auto
MME-Realworld [Fu et al., 2023] Holistic Evaluation of Perception and Cognition P R G Auto
MEGA-BENCH [Chen et al., 2024b] Large-scale suite with diverse real-world tasks P R G G Hybrid
Table 6: Further benchmarks assessing the “Thinking with Images” paradigm across Perception Reasoning, Code Generation, Chart Reasoning, and other Real-world Applications. The columns detail each benchmark’s Characteristics; the core cognitive Scenarios tested (P: Active Perception, R: Visual Reasoning, G: Generative Planning); the Eval Setup (M: multi-choice, G: generation); and the evaluation Metrics (Auto, LLM, Hybrid).

### v3

@@ 8 Future Directions
Although “Thinking with Images” marks a paradigm shift in artificial intelligence, we are still at the early stages of realizing its full potential. Looking forward, this section explores the pivotal research avenues that will propel the field into its next era, focusing on the innovations required to unlock more efficient, secure, and profoundly capable forms of visual cognition.
@@ 8.1 Computational and Cognitive Efficiency
A major challenge for “Thinking with Images” is its high computational cost (Ning et al., 2024; Zhou et al., 2024c). Current methods often rely on a sequence of visual steps, such as calling an external tool or generating an intermediate image. Each step adds significant delay and requires a lot of computation, making these models too slow for real-time applications (Liu et al., 2024b). Future research should find ways to achieve powerful visual reasoning without costly execution. This could involve teaching models to compress a long reasoning process into fewer steps, or to predict the final outcome without needing to render every single intermediate image (Lee et al., 2025).
A deeper challenge is improving cognitive efficiency, which is about using the right amount of effort for a given task (Qu et al., 2025; Huang et al., 2025c). A truly smart system should not use its most powerful and expensive thinking abilities for every simple problem (Wang et al., 2025n). The next generation of models must learn to decide when a quick textual answer is enough, and when a more careful visual analysis is actually needed. This is like a human deciding whether to solve a math problem in their head or to pull out a piece of paper and a pencil. Future work could train models with reward systems that penalize unnecessary visual steps, encouraging them to be more resourceful. The goal is to move from models that follow a fixed procedure to models that can flexibly switch between fast, simple thinking and slow, deliberate visual reasoning.
Achieving this dual efficiency will also require new model architectures (Deng et al., 2025a; Chen et al., 2025b). Instead of calling slow external tools, future models could have small, built-in modules designed for common visual tasks, like searching for an object or comparing two elements. These specialized modules could work directly on the model’s internal data representations, or latent space, avoiding the slow process of creating and analyzing full pixels. The objective is to develop models that can think with images while also demonstrating a sense of purpose and economy, making advanced visual reasoning a scalable and accessible technology.
@@ 8.2 Safe and Trustworthy Visual Cognition
The ability to “Think with Images” creates new safety and ethical challenges (Andriushchenko et al., 2024; Qu et al., 2024b). The most direct danger is the potential for misuse. Because these models can generate a series of images that seem to form a logical argument, a malicious user could guide a model to create a visual story that supports a false claim (Nguyen et al., 2022). This goes beyond a single deepfake. It enables the automated creation of entire disinformation campaigns, where fabricated visual evidence is presented step-by-step to make a lie seem true. The challenge, therefore, is not just blocking harmful images, but ensuring the entire thinking process cannot be hijacked to build a misleading narrative (Bai et al., 2022).
The security of these models is also a major concern due to their complex structure. The points of attack are no longer just the initial inputs, but the entire reasoning loop (Liu et al., 2024c, e, f). An attacker could poison the external tools a model relies on, corrupting its analysis from the start. A more subtle attack could involve designing an input that tricks the model into making a bad choice, such as focusing on a distracting detail or selecting the wrong tool for the job. In this way, the threat moves from simply causing a wrong output to corrupting the very process of thinking itself. This requires new security measures to protect a model’s decisions, tools, and internal imagination.
Finally, the thinking process of these models is often a "black box", which makes it hard to trust them. Hidden biases can affect not just the final output, but also the internal reasoning steps (Howard et al., 2024). For example, a model might learn to be more critical or generate stereotypical images only when reasoning about certain groups of people. This could create a hidden, self-reinforcing cycle of prejudice that is very difficult to spot. Making these systems trustworthy requires more than just explaining a result (Aflalo et al., 2022). We need new methods to see and review the model’s internal conversation of tool use and image generation. Without this transparency into how a model thinks with images, we cannot reliably find and fix internal biases, which prevents us from building models whose reasoning is not only powerful but also provably fair (Chang et al., 2024).
@@ 8.3 Novel Benchmarks and Evaluation Methodologies
A key problem for the “Thinking with Images” paradigm is that current benchmarks are inadequate (Li et al., 2024e). Most datasets only check the final answer and do not examine the reasoning process. This creates a loophole: a model can get the right answer for the wrong reasons, for example by using shortcuts instead of genuinely thinking with the image (Chen et al., 2024d). To address this, future work should focus on two main areas: creating more complex scenarios that require thinking with images, and developing new ways to evaluate the thinking process itself.
The first area of future work is to build new benchmarks with tasks that cannot be solved without actively thinking with images. One important direction is to design interactive and multi-step problems. For example, a task could require a model to solve a virtual puzzle, where each action changes the visual scene and reveals the next clue. Another promising direction is to create a greater variety of meaningful and challenging tasks that require visual construction. This means the model must generate a new visual element to find the solution, such as drawing an auxiliary line on a geometry diagram or sketching a simple machine to solve a physics problem (Wu et al., 2024h). In these scenarios, simply looking at the image once is not enough; success depends on a continuous process of visual interaction and creation.
The second area of future work is to develop new methods for evaluating the thinking process, not just the final answer (Song et al., 2025). Instead of only using accuracy, future metrics must assess the quality of the “Thinking with Images” chain. For example, we need evaluation methods that can check if each step of a model’s thought is grounded in the visual evidence, penalizing any steps that are made up. We also need to measure the logical coherence and necessity of the process. A practical way to test necessity is to remove an intermediate visual step generated by the model and see if it can still arrive at the correct solution. Establishing these process-oriented evaluation frameworks is a foundational step toward building more transparent and trustworthy multimodal AI systems.
@@ 8.4 Thinking with Audio, Video, and the World
The ultimate frontier of this research results in “Thinking with the world”, where the paradigm shifts from detached observation to embodied agency (Brohan et al., 2023). Here, reasoning becomes an integral component of a continuous perception-action loop. The model’s thoughts directly inform its physical actions, and the sensory consequences provide immediate feedback that grounds its understanding (Fu et al., 2024b). Its generative power evolves into a mechanism for mental rehearsal, allowing the agent to simulate action sequences before committing to a physical choice. This process grounds abstract concepts in tangible reality, defining an object not by its appearance but by its physical affordances in a planned task (Liu et al., 2024g).
@@ 8.5 Open Questions and Outlook
Rather than being mutually exclusive, the “Thinking with Images” paradigm and an agent framework share a interconnected and complementary relationship. The most effective way to conceptualize this is to view the “Thinking with Images” paradigm as the cognitive engine within the agent’s machine. An agent provides the essential architecture for worldly interaction, equipped with perception sensors, action effectors, and memory systems. The “Thinking with Images” paradigm provides the advanced reasoning power that makes the agent’s actions intelligent rather than merely reactive. This synergy becomes tangible when we consider how the stages outlined in this survey empower an agent’s capabilities. An agent leverages Stage 1 by learning to orchestrate external visual tools. It evolves by using Stage 2 to programmatically analyze visual data. Ultimately, a truly advanced agent embodies Stage 3, using intrinsic imagination to simulate the consequences of potential actions. Therefore, the research presented in this survey is not an alternative to agent-based AI but is instead a foundational step toward building autonomous systems that possess a deeper, more human-like capacity for visual understanding and intelligent action.
@@ 9 Conclusion
The journey towards true visual cognition is still in its early stages. Significant challenges in efficiency, robustness, and generalization must still be overcome. The path forward requires developing more well-developed world models, exploring abstract visual representations beyond pixels, and building unified frameworks for human-AI collaboration. Ultimately, the research presented here marks a foundational step toward creating AI agents that not only see the world, but reason within it as well, equipped with a mind’s eye that is integral to their intelligence.
Contributions
The contributions of all authors are listed as follows: Zhaochen Su drafted the abstract, introduction (§1), foundations (§2), and the introductory sections for each stage (§3.1, §4.1-4.2, §5.1-5.2). For the core methods, authorship was divided by approach: Zhenhua Liu drafted the prompting sections (§3.3.1, §4.4.1), Peng Xia the SFT sections (§3.3.2, §4.4.2, §5.4.1), and Yan Ma the RL sections (§3.3.3, §4.4.3, §5.4.2). Additionally, they authored key framing sections for their respective stages: Zhenhua for Stage 1 (§3.2, §3.4), Peng for Stage 2 (§4.3, §4.5), and Yan for Stage 3 (§5.3, §5.5). Moreover, Hangyu Guo drafted §6, and Jiaqi Liu drafted §7. Zhaochen Su and Xiaoye Qu drafted §8.
For visual elements, Zhaochen Su created the abstract figure and Figures 2, 3, and 5. Figure 1 was contributed by Peng Xia, Zhenhua Liu, and Kaide Zeng, while Figure 4 was contributed by Hangyu Guo. Yanshu Li provided visual assets and initial drafts. All tables (Tables 1, 2, 3) were a collaborative effort by Zhaochen, Zhenhua, Peng, Yan, and Kaide.
Junxian He and Yi R. (May) Fung supervised the project, providing invaluable guidance on its overall direction, structure, and refinement. Zhengyuan Yang, Linjie Li, Yu Cheng, and Heng Ji provide insightful feedback and critical suggestions on the manuscript.